OpenStudy (anonymous):

Solve the following systems of equations. x^2+y^2=16 x+y=4

OpenStudy (kittiwitti1):

Isolate one of the variables in one equation, then substitute the "isolated variable" equation you get into the other original equation. example:\[x=-y+4\]\[(-y+4)^{2}+y^{2}=16\]\[\text{solve for }y\]

OpenStudy (kittiwitti1):

After that, it becomes easier to solve for the other variable. :)

OpenStudy (anonymous):

where did you get y^2

OpenStudy (kittiwitti1):

It's part of one of the original equations.\[x^{2}+y^{2}=16\text{, if you isolate x in the other equation, you can replace the }x\text{ in this equation}\]

OpenStudy (kittiwitti1):

*you can replace the x in this equation

OpenStudy (kittiwitti1):

Basically, the substitution method, you know what I mean?

OpenStudy (anonymous):

yes

OpenStudy (kittiwitti1):

Okay, so do you know where to go from here?

OpenStudy (anonymous):

y=0

OpenStudy (kittiwitti1):

o-o You solved and got that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=4 y=0

OpenStudy (kittiwitti1):

well x=4 y=0 is a valid result but the answer can also be x=0 y=4

OpenStudy (kittiwitti1):

Since you can isolate y instead of x which would cause the second set of answers if you want whole #'s that is because you could also do:\[\sqrt{8}^{2}+\sqrt{8}^{2}=16\]